I am to recall that, when interest is compounded continuously, the balance in an account after t years is given by
A = Pe^rt,
where P is the initial investment and r is the interest rate.
Suppose that $2000 is invested at a rate of 6% per year compounded continuously. What is the balance after 1 yr? After 2 yrs?
Thank you in advance.

Answer by stanbon(48535)

A = Pe^rt,
where P is the initial investment and r is the interest rate.
Suppose that $2000 is invested at a rate of 6% per year compounded continuously. What is the balance after 1 yr? After 2 yrs?
-------------------
A(1) = 2000*e^(0.06*1)
A(1) = $2123.67
----------------
A(2) = 2000*e^(0.06*2)
A(2) = $2254.99
====================
Cheers,
Stan H.